# Are formal and functional linguistic mechanisms dissociated in language models?

### v3

@@ Abstract
Although large language models (LLMs) are increasingly capable, these capabilities are unevenly distributed: they excel at formal linguistic tasks, such as producing fluent, grammatical text, but struggle more with functional linguistic tasks like reasoning and consistent fact retrieval. Inspired by neuroscience, recent work suggests that to succeed on both formal and functional linguistic tasks, LLMs should use different mechanisms for each; such localization could either be built-in or emerge spontaneously through training. In this paper, we ask: do current models, with fast-improving functional linguistic abilities, exhibit distinct localization of formal and functional linguistic mechanisms? We answer this by finding and comparing the “circuits”, or minimal computational subgraphs, responsible for various formal and functional tasks. Comparing 5 LLMs across 10 distinct tasks, we find that while there is indeed little overlap between circuits for formal and functional tasks, there is also little overlap between formal linguistic tasks, as exists in the human brain. Thus, a single formal linguistic network, unified and distinct from functional task circuits, remains elusive. However, in terms of cross-task faithfulness—the ability of one circuit to solve another’s task—we observe a separation between formal and functional mechanisms, suggesting that shared mechanisms between formal tasks may exist.
@@ 1 Introduction
Much work has attempted to ameliorate these problems via retrieval-augmented generation or the use of chain-of-thought reasoning (Gao et al., 2024; Wei et al., 2022). However, Mahowald et al. offer another solution: perhaps LLMs would have stronger functional linguistic abilities if formal and functional linguistic abilities were as distinct in LLMs as they are in the human brain. Such a dissociation in LLMs could come about in two ways: either it could be explicitly built into them, or it could arise naturally via training, or the model’s inductive biases. Today’s LLMs have no explicit formal-functional modularity, but whether any has arisen despite that fact is an open question.
@@
To investigate whether formal and functional competence are distinctly localized within the LLMs, we draw on techniques from LLM interpretability, and in particular, mechanistic interpretability, which studies the mechanisms that underlie LLMs’ behavior using low-level causal methods (Ferrando et al., 2024). Concretely, we study the formal-functional distinction using circuits: a circuit is a small path through the LLM that contains all of the mechanisms underlying its behavior on a task of interest (Olah et al., 2020; Wang et al., 2023). Circuits provide causal guarantees that the localization found is correct, as all parts of the model outside of the circuit can be ablated without changing model behavior; see Section 3 for more details. This targeted, causal evidence for the correctness of our localization is a notable advantage of our framework, as such evidence is hard to come by in human brains. In humans, one must rely on either natural experiments (e.g., individuals with brain damage) or coarse-grained experiments using transcranial magnetic stimulation; optogenetics allows for finer-grained causal manipulation of neurons, but only in non-human subjects.
We thus translate Mahowald et al.’s hypothesis about emergent dissociation between formal and functional linguistic abilities in LLMs into the LLM circuit analysis framework. First, we identified 5 tasks involving formal linguistic competence, and 5 tasks involving functional linguistic competence. We next selected 5 LLMs, and found the circuits responsible for their behavior in these tasks. We then measured the similarity between each pair of task circuits, focusing in particular on similarities both within and across the formal and functional task groups; see Figure 1 for an overview.
@@
In summary, we find evidence for emerging formal-functional dissociation in LLMs, though only in terms of cross-task faithfulness. That is, formal task circuits perform other formal tasks better than they perform functional ones; functional tasks similarly struggle to perform formal tasks. However, our overlap studies suggest against the existence of one unified region for all language tasks. In performing these analyses, we apply mechanistic interpretability techniques to a question from neuroscience in LLMs for the first time. We moreover conduct a study of circuits across LLMs, employing both the greatest number and largest size LLMs of such a study to date. We release the code for our experiments, as well as the efficient circuit-finding tools within them, at https://github.com/hannamw/formal-functional-dissociation.
@@ 2.1 The Language Network in the Human Brain
The bulk of these findings come from experiments using a relatively simple setup used to localize brain regions responsible for a given competence. In such studies, brain activations (typically measured using functional magnetic resonance imaging; (fMRI) are measured in response to two contrasting stimuli (or while performing two contrasting tasks), where only one is relevant to the phenomenon of interest. Areas that differentially respond to the phenomenon of interest are inferred to be involved in its processing. For example, Fedorenko et al. (2010) localize the language network by finding regions that activate on well-formed sentences as opposed to lists of non-words.
We note that the aforementioned views on the language network are somewhat controversial. For example, Murphy and Woolnough (2024) criticize this body of work for its (over-)reliance on fMRI selectivity evidence, while others question the idea of separating formal and functional language use at all (Forkel and Hagoort, 2024). In this paper, we make no claims about the formal language network in the brain. Rather, we seek to translate the Mahowald et al.’s claim into LLMs, and test whether it holds.
@@ 2.2 Formal and Functional Linguistic Competence in Language Models
Mahowald et al. divide formal and functional linguistic competence into subdomains (Table 1). Formal language competence includes subdomains like phonology, morphology, syntax, and lexical semantics. For Mahowald et al., phonology corresponds to the rules governing valid wordforms (i.e., phonotactics), while morphology involves the correct ordering of morphemes. Syntax involves not only correct word order, but also higher-level abilities like agreement (e.g. between subjects and verbs). Lexical semantics entails using words correctly according to their part of speech, lexical category, or meaning. Mahowald et al. distinguish this category from semantics more broadly: general conceptual knowledge belongs to functional language competence.
@@ 3.1 Definitions
A model’s circuit for a given task is the minimal computational subgraph of the model that is faithful to its behavior on the task (Wang et al., 2023; Hanna, Liu, and Variengien, 2023). That is, even if all parts of the model outside of the circuit are ablated (or instead corrupted), model behavior will not change. See Hanna, Pezzelle, and Belinkov (2024) or Miller, Chughtai, and Saunders (2024) for reviews of circuits literature.
@@
In this study, we focus on LLMs using the transformer architecture (Vaswani et al., 2017). Such a model’s computational graph describes the computations it performs, and can be viewed as a directed graph that begins at the model’s inputs, flows through its components, i.e. attention heads and multi-layer perceptrons (MLPs), and ends at its logits (Conmy et al., 2023; Hanna, Liu, and Variengien, 2023). In most autoregressive transformers, each component’s input is directly added to its output via a residual connection; as a result, each component’s input is the sum of the outputs of previous components. Thus, in our computational graph, every component has an edge to components that come later in the model, and the input to a given node $v$ is the sum of the output of all nodes $u$ with an edge to $v$. A circuit should identify those edges that are important for the model’s ability to perform the task at hand. See Figure 2 for a toy example of a circuit in a 2-layer transformer model’s computational graph.
@@
A task in circuits analysis consists of clean and corrupted inputs, expected outputs, and a metric to measure task performance. For example, a subject-verb agreement task (like that of Linzen, Dupoux, and Goldberg, 2016; Newman et al., 2021) might consist of clean inputs like $s=$ The keys on the cabinet; the expected output would be a plural-conjugated verb, like are. The corrupted outputs are drawn from the same task distribution, but elicit very different behavior; here, the corrupted input might be $s^{\prime}=$The key on the cabinet, which elicits a singular-conjugated verb like is.
The metric could then be the difference in probabilities assigned to plural and singular verbs, given $s$; i.e., $m=p(\text{plural-verb}|s)-p(\text{singular-verb}|s)$. Note that the metric should be continuous, both for use with circuit-finding methods (see Section 3.2) and to allow for observation of incremental changes in model behavior as changes are made to the circuit. Tasks should be solvable by the model in which the circuit is to be found; if the model cannot solve the task, there may not be any task behavior to localize.
@@
A circuit is faithful if the model’s task behavior stays the same, even when all nodes and edges outside the circuit are ablated. At a high level, this entails running our model on clean inputs $s$, but replacing the activations of the nodes and edges outside the circuit, with activations taken from corrupted inputs $s^{\prime}$. As the vast majority of the model is outside the circuit, the model should behave as if it were being run on $s^{\prime}$—unless we have correctly localized all task-relevant nodes and edges.
More formally, and following Hanna, Pezzelle, and Belinkov (2024), we test model faithfulness by running the model on $s$ and performing the following causal intervention. Let $v$ be a node in the whole model’s computational graph $G=(V,E)$. Let $z_{u}$ denote the activation of a given node $u$ during the current forward pass, and $z^{\prime}_{u}$ denote its activation on corrupted inputs $s^{\prime}$. Let $C=(V_{C},V_{E})$ denote our circuit. For each non-input node $v$, we set its input to
@@
Having done this, we measure $m_{C}$, the circuit’s performance. Denote by $m_{\emptyset}$ the model’s performance when entirely corrupted (i.e., when it is run on the corrupted input). We can then compute a measure of normalized faithfulness (Marks et al., 2024):
@@ 3.2 Finding Circuits
[…]
@@
where $(z_{u}^{\prime}-z_{u})$ indicates the change in $u$’s activation upon corruption, and $\nabla_{v_{in}}m(s)$ is the change in metric when $v$’s input changes. Computing the activations $z_{u}^{\prime}$ and $z_{u}$ of all nodes $u$ requires two forward passes, while computing $\nabla_{v_{in}}m(s)$ for all nodes $v$ requires one backward pass. We can thus compute $\hat{\text{IE}}$ for all edges in constant time.
As EAP is often inaccurate in practice, we instead use Hanna, Pezzelle, and Belinkov’s (2024) EAP with integrated gradients (EAP-IG), which improves upon EAP by computing $\nabla_{v_{in}}m$ at intermediate points between $s$ and $s^{\prime}$. Having thus computed each edge’s IE, we find the circuit by taking the top-$n$ edges by $|\hat{\text{IE}}|$. We take edges by the absolute value of $\hat{\text{IE}}$ to find a circuit that is a complete explanation of the model’s task behavior, containing all relevant nodes and edges, even if they harm performance.
@@ 3.3 Why use circuits to localize formal and functional linguistic competence?
These issues have played an important role in prior work attempting to localize a language network in LLMs. Zhang et al. (2024), for example, localize a “core linguistic region” consisting of model parameters that harm model abilities cross-lingually when set to zero. However, such zero ablations can cause harm unrelated to the importance of the units ablated (see Section 2.3); moreover, they do not check that this zero-ablation harms language selectively. Thus, we cannot be sure whether this network is crucial for LLMs’ language abilities, or their abilities in general.
@@ 4 Tasks and Data
In our main experiments, we consider 10 tasks that give us broad coverage over most of the subdomains of formal and functional linguistic competence described by Mahowald et al.. We exclude some such categories (like phonology) as they are impossible to test in text-based LMs. For others (social reasoning), LM abilities are generally poor, while circuits are best found for tasks models perform well; LM are quite competent on all tasks we study here (see App. A for details). See Table 2 for an overview of tasks. While we introduce some tasks that are new to circuit analysis, most have previously been studied in the circuits literature. We note, however, that because prior work studied these tasks in smaller or older models, it is difficult to make direct comparisons between the circuits we find, and the circuits prior work found.
@@
Formal Subject-Verb Agreement (SVA) Syntax The keys on the cabinet [are]
@@
Functional Indirect Object Identification (IOI) Situation Modeling Alice and Bob went to the store. Then Alice gave a bottle of water to [Bob]
@@ 4.1 Formal Tasks
Subject-Verb Agreement (SVA) is a Syntactic task that gives models inputs like $s=$ “The keys to the cabinet”, and expects verbs that agree with the subject keys. Its corrupted variant inverts the plurality of the subject; here keys would change to key. We measure task performance as the probability assigned to verbs that agree with the subject, minus that assigned to verbs that do not. For this task, we adapted Newman et al.’s (2021) SVA data.
@@
Gendered-Pronoun Agreement is a Syntactic task that gives models inputs involving explicitly gendered entities, like $s=$ “The heroine went home because”. The expected output is the corresponding gendered pronoun, she. Its corrupted variant replaces the subject with the corresponding opposite-gender noun (here, hero); task performance is measured as ${logit}(\text{she})-{logit}(\text{he})$. Vig et al. (2020) identified neurons causally responsible for models’ ability to perform this task in a gender-biased scenario; we adapt their data, adding explicitly gendered entities.
@@
Negative Polarity Item (NPI) usage is a Syntactic task that gives models inputs that do or do not license NPIs, like ever or any. For example, the input $s=$ “The customer that the managers liked has” could be continued by “never”, but not by the NPI “ever”. In each corrupted variant, we add an NPI-licensing word to $s$, or remove it if it already exists; in our example, $s^{\prime}$ could be “No customer that the managers liked has”, which licenses the use of the NPI ever. Task performance is measured as ${logit}(\text{never})-{logit}(\text{ever})$. We adapted this task’s data from the corresponding SyntaxGym task (Gauthier et al., 2020).
Hypernymy is a Lexical Semantic task that gives models inputs like $s=$ “Roses are a type of”, and expects outputs like flower. Its corrupted variant replaces the hyponym (roses) with that of another type (e.g. diamonds); the resulting metric is $p(\text{flower}|s)-p(\text{gem}|s)$. We use the task data from Hanna, Pezzelle, and Belinkov (2024).
@@
The Wug Test is a Morphological/Syntactic task that tests models’ abilities to generate singular and plural forms of nonce words (Berko, 1958). It gives models inputs containing nonce words like $s=$ “First, there was one trult. Now there are two”, and expects outputs like trults. Its corrupted variant reverses the number of entities present, as in $s^{\prime}=$ “First, there were two trults. Now there is one”. Task performance is measured as $p(\text{trults})-p(\text{trult})$. We generate new nonce words using Wuggy (Keuleers and Brysbaert, 2010) to avoid LMs having previous exposure. The Wug test requires morphological abilities (namely, to generate the plural or singular of a nonce word) but also requires syntactic abilities (to understand that the nonce word and preceding numeral must agree in number).

### v4

@@ Abstract
Although large language models (LLMs) are increasingly capable, these capabilities are unevenly distributed: they excel at formal linguistic tasks, such as producing fluent, grammatical text, but struggle more with functional linguistic tasks like reasoning and consistent fact retrieval. Inspired by neuroscience, recent work suggests that to succeed on both formal and functional linguistic tasks, LLMs should use different mechanisms for each; such localization could either be built-in or emerge spontaneously through training. In this paper, we ask: do current models, with fast-improving functional linguistic abilities, exhibit distinct localization of formal and functional linguistic mechanisms? We answer this by finding and comparing the “circuits”, or minimal computational subgraphs, responsible for various formal and functional tasks. Comparing 5 LLMs across 10 distinct tasks, we find that while there is indeed little overlap between circuits for formal and functional tasks, there is also little overlap between formal linguistic tasks, as exists in the human brain. Thus, a single formal linguistic network, unified and distinct from functional task circuits, remains elusive. However, in terms of cross-task faithfulness—the ability of one circuit to solve another’s task—we observe a separation between formal and functional mechanisms, with formal task circuits achieving higher performance on other formal tasks. This suggests the existence of a set of formal linguistic mechanisms that is shared across formal tasks, even if not all mechanisms are strictly necessary for all formal tasks.
@@ 1 Introduction
Much work has attempted to ameliorate these problems via retrieval-augmented generation or the use of chain-of-thought reasoning (Gao et al., 2024; Wei et al., 2022). However, Mahowald et al. offer another solution: perhaps LLMs would have stronger functional linguistic abilities if formal and functional linguistic abilities were as distinct in LLMs as they are in the human brain. Such a dissociation in LLMs could come about in two ways: either it could be explicitly built into them, or it could arise naturally via training, or the model’s inductive biases. Today’s LLMs have no explicit formal-functional modularity; however, it remains unknown whether formal-functional modularity has nonetheless arisen.
@@
To investigate whether formal and functional competence are distinctly localized within the LLMs, we draw on techniques from LLM interpretability, and in particular, mechanistic interpretability, which studies the mechanisms that underlie LLMs’ behavior using low-level causal methods (Ferrando et al., 2024). Concretely, we study the formal-functional distinction using circuits: a circuit is a small path through the LLM that contains all of the mechanisms underlying its behavior on a task of interest (Olah et al., 2020; Wang et al., 2023). Circuits provide causal guarantees that the localization found is correct, as all parts of the model outside of the circuit can be ablated without changing model behavior; see Section˜3 for more details. This targeted, causal evidence for the correctness of our localization is a notable advantage of our framework, as such evidence is harder to come by in human brains. [^1] In humans, one must rely on either natural experiments (e.g., individuals with brain damage) or less precise causal methodologies. These are numerous, ranging from more targeted methods such as transcranial magentic stimulation or focused ultrasound, to the more general interventions used in psychedelics research, but they are overall less precise than our causal interventions; optogenetics allows for finer-grained causal manipulation of neurons, but only in non-human subjects.
We thus translate Mahowald et al.’s hypothesis about emergent dissociation between formal and functional linguistic abilities in LLMs into the LLM circuit analysis framework. First, we identified 5 tasks involving formal linguistic competence, and 5 tasks involving functional linguistic competence. We next selected 5 LLMs, and found the circuits responsible for their behavior in these tasks. We then measured the similarity between each pair of task circuits, focusing in particular on similarities both within and across the formal and functional task groups; see Figure˜1 for an overview.
@@
In summary, we find evidence for emerging formal-functional dissociation in LLMs, though only in terms of cross-task faithfulness. That is, formal task circuits perform other formal tasks better than they perform functional ones; functional tasks circuits similarly struggle to perform formal tasks. However, our overlap studies suggest against the existence of one unified region necessary for all language tasks. Rather, there may be a broad pool of mechanisms shared across formal tasks; some formal tasks may be solved by multiple mechanisms (circuits) from this pool. In performing these analyses, we apply mechanistic interpretability techniques to a question from neuroscience in LLMs for the first time. We moreover conduct a study of circuits across LLMs, employing both the greatest number and largest size LLMs of such a study to date. We release the code for our experiments, as well as the efficient circuit-finding tools within them, at https://github.com/hannamw/formal-functional-dissociation.
@@ 2.1 The Language Network in the Human Brain
The bulk of these findings come from experiments using a relatively simple setup used to localize brain regions responsible for a given competence. In such studies, brain activations (typically measured using functional magnetic resonance imaging; (fMRI) are measured in response to two contrasting stimuli (or while performing two contrasting tasks), where only one is relevant to the phenomenon of interest. Areas that differentially respond to the phenomenon of interest are inferred to be involved in its processing. Using these methods, Fedorenko et al. (2010) localize the language network by finding regions that activate on well-formed sentences as opposed to lists of non-words. Note that the differential response requirement ensures that low-level regions that are activated by both contrasting stimuli are excluded from the localized region: for example, the language network excludes low-level speech processing areas, as these are activated by both sentences and non-word lists.
In summary, there is a wide body of evidence for the formal-function dissociation. Though this dissociation is not universally supported (Murphy and Woolnough, 2024; Forkel and Hagoort, 2024), the evidence in its favor is substantial, spanning over a decade of research, and consisting of studies using a wide variety of languages and non-linguistic tasks. We believe this body of neuroscientific evidence provides a solid theoretical motivation for our study on whether emergent formal-functional modularity, as suggested by Mahowald et al., exists in today’s LLMs. However, see Section˜9 for discussion of dissenting views, and how our results might be interpreted from alternative perspectives.
@@ 2.2 Formal and Functional Linguistic Competence in Language Models
Mahowald et al. divide formal and functional linguistic competence into subdomains (Table˜1). Formal language competence includes subdomains like phonology, morphology, syntax, and lexical semantics. For Mahowald et al., phonology corresponds to the rules governing valid wordforms (i.e., phonotactics), while morphology involves the correct ordering of morphemes. Syntax involves not only correct word order, but also higher-level abilities like agreement (e.g. between subjects and verbs). Lexical semantics entails using words correctly according to their part of speech, lexical category, or meaning. Mahowald et al. distinguish this category from semantics more broadly: general conceptual knowledge belongs to functional language competence.
@@ 3.1 Definitions
A model’s circuit for a given task is the minimal computational subgraph of the model that is faithful to its behavior on the task (Wang et al., 2023; Hanna, Liu, and Variengien, 2023). That is, even if all parts of the model outside of the circuit are ablated (or instead corrupted), model behavior will not change. See Hanna, Pezzelle, and Belinkov (2024) or Miller, Chughtai, and Saunders (2024) for reviews of the circuits literature.
@@
In this study, we focus on LLMs using the transformer architecture (Vaswani et al., 2017). Such a model’s computational graph describes the computations it performs, and can be viewed as a directed graph that begins at the model’s inputs, flows through its intermediate components, and ends at its logits (Conmy et al., 2023; Hanna, Liu, and Variengien, 2023). In this paper, the intermediate components we study are either individual attention heads or multi-layer perceptrons (MLPs); thus, every node in our graph is an input (token) node, attention head, MLP, or an output (logit) node.
Most autoregressive transformers can be conceptualized as having a residual stream. This residual stream is initialized with the input embeddings (and potentially positional embeddings); then, each component takes the stream as input, and adds its own output to this stream. As a result, each component’s input is the sum of the outputs of previous components. This means that each component has a direct effect (unmediated by other components) to every component downstream of it; it also has an indirect effect (mediated by other components). Edges in our computational graph represent these direct effects, so every component has an edge to all components that come later in the models. The input to a given node $v$ is the sum of the output of all nodes $u$ with an edge to $v$. A circuit should identify those edges that are important for the model’s ability to perform the task at hand. See Figure˜2 for a toy example of a circuit in a 2-layer transformer model’s computational graph.
@@
A task in circuits analysis consists of clean and corrupted inputs, expected outputs, and a metric to measure task performance. For example, a subject-verb agreement task (like that of Linzen, Dupoux, and Goldberg, 2016; Newman et al., 2021) might consist of clean inputs like $s=$ The keys on the cabinet; the expected output would be a plural-conjugated verb, like are. Each clean input is paired with a corrupted input, which is drawn from the same task distribution, but is crafted to elicit very different behavior from the model. Here, the corrupted input might be $s^{\prime}=$The key on the cabinet, which elicits a singular-conjugated verb like is. $s$ and $s^{\prime}$ should have the same length in tokens.
Our metric could then be the difference in probabilities assigned to correctly and incorrectly conjugated verbs, given $s$; i.e., $m=p(\text{plural-verb}|s)-p(\text{singular-verb}|s)$. Note that $m$ should be high when the model is run on $s$, but low (negative) when it is run on $s^{\prime}$, i.e. $s^{\prime}$ elicits the opposite model behavior compared to $s$ with respect to $m$.
The metric should be continuous, both for use with circuit-finding methods (see Section˜3.2) and to allow for observation of incremental changes in model behavior as changes are made to the circuit. Tasks should be solvable by the model in which the circuit is to be found; if the model cannot solve the task, there may not be any task behavior to localize. The model’s performance on a given task example should ideally be measurable in one forward pass.
@@
A circuit is faithful if the model’s task behavior stays the same, even when all nodes and edges outside the circuit are corrupted. At a high level, corruption entails running our model on clean inputs $s$, but replacing the activations of the nodes and edges outside the circuit, with activations taken from corrupted inputs $s^{\prime}$. As the vast majority of the model is outside the circuit, the model will behave as if it were being run on $s^{\prime}$, and the value of the metric will drop—unless we have correctly localized all task-relevant nodes and edges.
More formally, and following Hanna, Pezzelle, and Belinkov (2024), we test model faithfulness by running the model on $s$ and performing the following causal intervention. Let $v$ be a node in the whole model’s computational graph $G=(V,E)$. Let $z_{u}$ denote the activation of a given node $u$ during the current forward pass, and $z^{\prime}_{u}$ denote its activation on corrupted inputs $s^{\prime}$. Let $C=(V_{C},E_{C})$ denote our circuit.
For each non-input node $v$, we set its input to
@@
Having done this, we measure $m_{C}$, the circuit’s performance. Denote by $m$ the model’s original performance on the task, and by $m_{\emptyset}$ the model’s performance when entirely corrupted (i.e., when it is run on the corrupted input). We can then compute a measure of normalized faithfulness (Marks et al., 2024):
@@ 3.2 Finding Circuits
A naive approach to finding circuits is to compute how much the model’s performance decreases when each edge is corrupted; this is the edge’s indirect effect (IE; Pearl, 2001). Note that IE can be positive, which implies that the edge promotes model performance, or negative, implying that it harms it. To find a circuit of size $n$, one could simply take the $n$ edges with the highest $|\text{IE}|$; we take the absolute value because we want to capture all edges that affect model behavior, positively or negatively. [^2] Past work has found components and edges that actively and systematically work against model performance (Wang et al., 2023), which are a relevant part of our model’s mechanism. However, as testing an edge requires one forward pass, this approach takes $O(|E|)$ forward passes, and both $|E|$ and the cost of a forward pass increase with model size.
@@
where $(z_{u}^{\prime}-z_{u})$ indicates the change in the activation of $u$ upon corruption, and $\nabla_{v_{in}}m(s)$ is the change in metric when the input to $v$ changes. Computing the activations $z_{u}^{\prime}$ and $z_{u}$ of all nodes $u$ requires two forward passes, while computing $\nabla_{v_{in}}m(s)$ for all nodes $v$ requires one backward pass. We can thus compute $\hat{\text{IE}}$ for all edges in a constant number of forward and backward passes, although the cost of each pass grows with model size.
[…]
@@ 3.3 Why use circuits to localize formal and functional linguistic competence?
These issues have played an important role in prior work attempting to localize a language network in LLMs. Zhang et al. (2024), for example, localize a “core linguistic region” consisting of model parameters that harm model abilities cross-lingually when set to zero. However, such zero ablations can cause harm unrelated to the importance of the units ablated (see Section˜2.3); moreover, they do not check that this zero-ablation harms language selectively. Thus, we cannot be sure whether this network is crucial for LLMs’ language abilities, or their abilities in general.
@@ 4 Tasks and Data
In our main experiments, we consider 10 tasks that give us broad coverage over most of the subdomains of formal and functional linguistic competence described by Mahowald et al.. We exclude some such categories (like phonology) as they are impossible to test in text-based LMs. For others (social reasoning), LM abilities are generally poor, while circuits are best found for tasks models perform well; LM are quite competent on all tasks we study here (see App. A for details). See Table˜2 for an overview of tasks. While we introduce some tasks that are new to circuit analysis, most have previously been studied in the circuits literature. We note, however, that because prior work studied these tasks in smaller or older models, it is difficult to make direct comparisons between the circuits we find, and the circuits prior work found.
@@
Formal
Subject-Verb Agreement (SVA) Syntax The keys on the cabinet [are]
@@
Functional
Indirect Object Identification (IOI) Situation Modeling Alice and Bob went to the store. Then Alice gave a bottle of water to [Bob]
@@ 4.1 Formal Tasks
Subject-Verb Agreement (SVA) is a Syntactic task that gives models inputs like $s=$ “The keys to the cabinet”, and expects verbs that agree with the subject keys. Its corrupted variant inverts the plurality of the subject; here keys would change to key. We measure task performance as the probability assigned to verbs that agree with the subject, minus that assigned to verbs that do not. For this task, we adapted Newman et al.’s (2021) SVA data. SVA is a classic form of syntactic agreement, and has been used to study the language network in humans (Fedorenko et al., 2020).
@@
Gendered-Pronoun Agreement is a Syntactic task that gives models inputs involving explicitly gendered entities, like $s=$ “The heroine went home because”. The expected output is the corresponding gendered pronoun, she. Its corrupted variant replaces the subject with the corresponding opposite-gender noun (here, hero); task performance is measured as ${logit}(\text{she})-{logit}(\text{he})$. Vig et al. (2020) identified neurons causally responsible for models’ ability to perform this task in a gender-biased scenario; we adapt their data, adding explicitly gendered entities. Gendered-pronoun agreement is also a classic form of syntactic agreement, which has often been used to study the language network (Rodd et al., 2010; Fedorenko et al., 2020).
@@
Negative Polarity Item (NPI) usage is a Syntactic task that gives models inputs that do or do not license NPIs, like ever or any. For example, the input $s=$ “The customer that the managers liked has” could be continued by “never”, but not by the NPI “ever”. In each corrupted variant, we add an NPI-licensing word to $s$, or remove it if it already exists; in our example, $s^{\prime}$ could be “No customer that the managers liked has”, which licenses the use of the NPI ever. Task performance is measured as ${logit}(\text{never})-{logit}(\text{ever})$. We adapted this task’s data from the corresponding SyntaxGym task (Gauthier et al., 2020). For a discussion of NPI in the context of formal semantics in the brain, see Panizza (2012).
Hypernymy is a Lexical Semantic task that gives models inputs like $s=$ “Roses are a type of”, and expects outputs like flower. Its corrupted variant replaces the hyponym (roses) with that of another type (e.g. diamonds); the resulting metric is $p(\text{flower}|s)-p(\text{gem}|s)$. We use the task data from Hanna, Pezzelle, and Belinkov (2024). We note that this task and its categorization as a formal (lexical semantic) task, rather than a functional (perhaps world knowledge) task, may be somewhat controversial. This classification stems from Mahowald et al.’s (2024) taxonomy, which specifies that lexical (and compositional) semantics are formal linguistic competences; other aspects of semantics fall under world knowledge. Hypernymy is a classic example of a lexical semantic relation, along with hyponymy, synonymy, and antonymy (Cruse, 1986); we thus consider it to be a lexical semantic and formal task. While past work in neurolinguistics has not studied hypernymy specifically, Fedorenko et al. (2020) operationalized lexical semantics via one of its sister relations, synonymy. In their study, participants determined if two sentences’ meanings were the same, when one word was replaced by its synonym; they observed that the language network responded during this task, providing evidence for synonymy being a formal task.
@@
The Wug Test is a Morphological / Syntactic task that tests models’ abilities to generate singular and plural forms of nonce words (Berko, 1958). It gives models inputs containing nonce words like $s=$ “First, there was one trult. Now there are two”, and expects outputs like trults. Its corrupted variant reverses the number of entities present, as in $s^{\prime}=$ “First, there were two trults. Now there is one”. Task performance is measured as $p(\text{trults})-p(\text{trult})$. We generate new nonce words using Wuggy (Keuleers and Brysbaert, 2010) to avoid LMs having previous exposure. The Wug test requires morphological abilities (namely, to generate the plural or singular of a nonce word) but also requires syntactic abilities (to understand that the nonce word and preceding numeral must agree in number). Past work has studied morphology’s role in the language network, though it studied verbal, rather than nominal morphology (Bozic et al., 2010).